What is the tenth term of the sequence
64, 16, 4, ...?

Found 2 solutions by Fombitz, richwmiller:
Answer by Fombitz(32388)
Answer by richwmiller(17219)

Tn = t1 * r^(n - 1)
T10 = 64 * 1/4^(9)
T10 = 64 * 1/4^(9)
T10 = 1/4^-3 * 1/4^(9)
T10=1/4^6
T10=1/4096